Jane is 26 years older than Eve, who is 37 years younger than Ray. In how many years will Jane be the same age as Ray is now?

Answer by ankor@dixie-net.com(22740)

Jane is 26 years older than Eve,
j = e + 26
who is 37 years younger than Ray.
e = r - 37
or
r = e + 37
In how many years will Jane be the same age as Ray is now?
let y = no. years for this to be true
j + y = r
replace j with (e+26); replace r with (e+37)
e+26 + y = e + 37
subtract e from both sides
26 + y = 37
y = 37 - 26
y = 11 years, Jane will be the same age as Ray is now
: